OpenStudy (anonymous):

Let f(x) = x2 - 16. Find f-1(x).

OpenStudy (sheraz12345):

Do u know how to take inverse

OpenStudy (anonymous):

Not really.

OpenStudy (anonymous):

\(f(x)=y\\y=x^2-16\\x=y^2-16\) now, solve for y.

OpenStudy (anonymous):

16x=y^2/16x

OpenStudy (anonymous):

nope. transpose 16 to the left side by adding 16 on both sides. Then you have to get the square root of both sides.